Determine whether the ordered pair is a solution of the given equation
(5,-5)
y=x-10

Answer by stanbon(75887)

Determine whether the ordered pair is a solution of the given equation
(5,-5)
y=x-10
--------
Substitute for x and for y to see if you get a true statement.
-5 = 5-10
-5 = -5
-------------
true, so the point is a solution of the given equation.
==================
Cheers,
Stan H.